14. What volume of gas will be produced when 35.0 g of calcium reacts with excess hydrochloric acid at SATP?


1
Expert's answer
2021-06-14T13:37:05-0400

The chemical reaction between calcium and hydrochloric acid is as follows:

Ca + 2HCl = CaCl2 + H2

From here, the number of moles of calcium equals:

n(Ca) = m(Ca) / Mr(Ca) = 35.0 g / 40 g/mol = 0.875 mol

According to the reaction, the number of moles of calcium equals the number of moles of hydrogen produced. As a result, n(H2) = 0.875 mol

Finally, the volume of gas is:

V(H2) = n(H2) × 22.41 L/mol = 0.875 mol × 22.41 L/mol = 19.6 L


Answer: 19.6 L